# Optimistic critics can empower small actors

### v1

@@ 1 Introduction
The paper is organized as follows: in section 2, we lay the groundwork and explain our experimental setup. In section 3, we show the performance effects when naively reducing a smaller actor across a variety of state-based and image-based continuous control tasks, and analyze what could be the cause of performance differences. In section 4, we focus on interventions that can gain back performance, specifically focusing on bias correction and value function underestimation. Finally, we conclude with discussions and avenues for future work in section 5.
@@ 2 Preliminaries
In eq. 2, $V_{\bar{\phi}}$ is the value function computed from $Q_{\phi}$ via $V_{\bar{\phi}}(x)=\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid x)}\left[Q_{\bar{\phi}}(x,a)-\alpha\log\pi_{\theta}(a\mid x)\right]$, where $\bar{\phi}$ are delayed target parameters (Mnih et al., 2015), and $\alpha$ is a learned Lagrange multiplier (we exclude its parameterization for simplicity of exposition). In their practical implementation, Haarnoja et al. (2018b) use two $Q$ value estimates with parameters $\phi_{1}$ and $\phi_{2}$, trained independently, and take their minimum in the update terms in equations 2 and 3, resulting in the following updated losses, with $V_{\bar{\phi}}(x)=\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid x)}\left[\min_{i\in\{1,2\}}Q_{\bar{\phi_{i}}}(x,a)-\alpha\log\pi_{\theta}(a\mid x)\right]$:
@@
It is important to note the interconnectedness of these losses: the actor influences the critic via the (soft) value function $V_{\bar{\phi}}$ used in eqs. 2 and 4, while the critic influences the actor via $Q_{\phi}$ in eqs. 3 and 5. Additionally, the actor influences the training dynamics of both given that it is in charge of data collection. Finally, note the use of the TD-error term in eqs. 2 and 4.
@@ 3 The impact of small actors
In the top row of fig. 1 we evaluate the impact on performance when reducing the size of these layers and can see a clear degradation in performance across all environments. We additionally measure $o_{\phi}$ on the critics, introduced by Nauman et al. (2024) as a measure of overfitting, defined as $o_{\phi}:=\frac{\mathbb{E}_{\mathcal{D}_{V}}TD_{\phi}}{\mathbb{E}_{\mathcal{D}}TD_{\phi}}$. Here, $\mathcal{D}_{V}$ is a validation dataset of size $11,000$, containing data sampled from a training run with a regular unmodified SAC agent, trained with a different random seed, and $TD_{\phi}$ is the temporal difference error. Higher values of $o_{\phi}$ are indicative of overfitting which, as seen in the middle row of fig. 1, are inversely correlated with the size of the actor. Finally, we report the fraction of dormant neurons, defined as the proportion of neurons that are 00 for every data point in the validation buffer, where higher levels of dormancy is associated with a loss of plasticity (Sokar et al., 2023; Lyle et al., 2024; Klein et al., 2024). In the bottom row of fig. 1 we see that the fraction of dormant neurons tends to be inversely correlated with actor size and performance, particularly for the environments where the performance loss is greatest, although to a lesser extent than $o_{\phi}$.
@@ 3.1 Smaller actors collect worse data
Training overparameterized neural networks on small datasets is a common cause for overfitting. Given that the actor is in charge of data collection and reducing its expressivity results in overfit critics, we continue our investigation by evaluating the quality of the data gathered by differently-sized actors. For this, we evaluate training on data collected by separate, and differently-sized, actors. Specifically, we train the regularly-sized actor $\pi_{r}$ with data provided by one of the smaller actors, where the data collection exactly mimics that obtained by the smaller actor during training. This is depicted by the black bars in the top row of fig. 2, where we can see the performance to be clearly correlated with actor size.
Nikishin et al. (2022) demonstrated the tendency of RL agents to overfit to early experience, affecting their plasticity and downstream performance. It is thus worth considering whether the quality of the training data on an actor is most important in the early stages of training. To evaluate this, in the bottom row of fig. 2 we explore training the smallest actor ($\pi_{xs}$) on data provided by the $\pi_{r}$ actor, again matching the data collection of the smaller actor. Our analyses here explore using the data from $\pi_{r}$ for only a fraction of training, and then switching to data collected by $\pi_{xs}$ itself. As more data is collected from the bigger actor $\pi_{r}$, the performance of $\pi_{xs}$ generally improves. We note that using all of the data from $\pi_{r}$ (i.e. at 100%) sometimes results in degraded performance; we hypothesize that this may be due to the tandem effect observed by Ostrovski et al. (2021). Overall, we see an improvement in performance in the environments most impacted by reducing the size of the actor for SAC. With DrQ, we do not see a pronounced effect when training the smallest actor on data from a regular-sized actor (see fig. 9), but similarly, a small trend may be observed for environments with the biggest degradation in performance with reduced actor sizes.
@@ 3.2 Smaller actors result in critic underestimation
The reduction in the quality of data gathered by small actors can possibly be attributed to under-exploration of the state space. This can often be a consequence of an overly-conservative critic which under-estimates values, as well as a low-entropy actor with low diversity in action selection. In fig. 3 we compare the average critic validation $Q$-values (computed on the same validation dataset) as well as the entropy of the actor’s action distribution $\pi$ of the smaller actors relative to the regularly-sized actor, and confirm that smaller actors result in $Q$-value underestimation, as well as a reduction in entropy during training (see fig. 11 for the same comparison throughout training). The observed underestimation is interesting, given that it stands in contrast to the over-estimation that’s commonly addressed in many popular algorithms (Hasselt, 2010; Hasselt et al., 2016; Fujimoto et al., 2018).
@@ 4.1 Average and maximal critics
We begin by a simple modification to the original SAC losses to directly address the observed value underestimation. Specifically, we replace the minimization of the two independent $Q$ estimates in equations 4 and 5 with either their mean ($\textrm{avg}(Q_{\phi_{1}},Q_{\phi_{2}})$) or their maximum ($\max(Q_{\phi_{1}},Q_{\phi_{2}}))$. As can be seen in the top and middle rows of fig. 4 and the top row of fig. 5, this approach can be quite effective at boosting the performance of small actors in SAC, sometimes even improving over the minimization approach with the regular sized model (e.g. hopper-hop). The bottom row of fig. 5 confirms that this technique does increase the validation value estimates. As can be seen in fig. 13, we find that the mean and the max approaches also improve several overfitting and plasticity metrics in the critics, most notably $o_{\phi}$ and the rank of the last hidden layer (Kumar et al., 2021; Nauman et al., 2024). However, they do not appear to have a notable impact on these metrics in the actor (see fig. 14). The results on the smallest actor on DrQ (bottom row of fig. 4) display a similar performance trend, although the results are less pronounced. We also observe a corresponding trend with an increase in validation Q values with the mean and max approaches in DrQ in fig. 10.
@@ 4.2 Critic regularization
Prior work has shown that resetting or regularizing the critic in particular improves plasticity (Ma et al., 2023; Nikishin et al., 2022; Liu et al., 2021) and can help mitigate value-estimation issues, particularly in the case of layer normalization, (Nauman et al., 2024), albeit with overestimation. Given both the increased overfitting observed in the critics, and how much value estimation is affected by smaller actors (see fig. 3), we investigate whether critic regularization alone can be effective mitigating this impact by applying a number of regularization techniques, focusing on SAC: (a) Layer Normalization (Ba et al., 2016); (b) Spectral Normalization (Miyato et al., 2018); (c) weight decay (van Laarhoven, 2017) with a regularization value of $0.01$ (Li et al., 2023a); (d) L2 distance from initialization (Kumar et al., 2024): with a value of $1\times 10^{-7}$ after tuning on the range $[5\times 10^{-8},1\times 10^{-4}]$ in increments of $0.5$ with quadruped-run; and (e) Network resets: resetting neural network layers during training (Nikishin et al., 2022). We apply layer normalization and spectral normalization to the second hidden layer in the critics, and we reset only the output layer of the critics every $50$K steps. Although many of these methods do appear to help with mitigating value under-estimation (bottom row of fig. 5), they do not appear to help much with performance (top row of fig. 5 and fig. 12). For DrQ, we investigate resetting the MLP of the critics (Nikishin et al., 2022) for the smallest actors, and similarly do not see a notable rescue effect (see fig. 10).
@@ 4.3 Addressing bias in the critic via actor representations
In asymmetric actor-critics methods, imbalances in information received by the critic versus the actor can lead to biased gradients that may negatively impact performance; Baisero & Amato (2022) and Lyu et al. (2022) propose to alleviate this by giving the (limited) information received by the actor as additional input into the critic. In our case, the critics do not receive privileged information over the actor, but we theorize that a similar effect may be occurring within the policy network due to potentially impacted information flow through the smaller actors. We attempt a similar bias correction by concatenating the latent state of the final hidden layer of the actor as input to the final hidden layer of the critics. The latent state is first projected through an untrained neural network layer to a size of $8$ to maintain consistency across actor sizes. As shown in the top row of fig. 5 and the bottom row of fig. 12, the bias correction performs similarly to other attempted critic regularization methods in SAC.
@@ 5 Discussion
Real-world problems are often subject to constraints such as latency, model size, and interpretability, which are largely absent in the academic benchmarks where machine learning solutions are developed. As such, it is imperative that we develop the necessary techniques for training reinforcement learning agents under such limitations. The use of small actors can help reduce latency, memory, and inference costs, and can help improve interpretability; these are all practical considerations, as ultimately it is a trained actor which will be deployed for action selection. Our work demonstrates that naïvely shrinking the actor can result in value underestimation, poor data collection, and ultimately degraded performance. We evaluated a number of approaches for mitigating this deterioration and found the most effective to be simply replacing the min operation with a mean or max when combining the values of the two critics (section 4.1).
It is often necessary to provide the actor with less information than the critic, as was employed by Vasco et al. (2024) to better match the inputs used by humans. In fig. 6 we explore whether this additional type of limitation on the actor may have a similar effect to what we observe when decreasing the size of the actor. To test this, we zero out two-thirds of the actor inputs in SAC (retaining every third dimension) and find that taking the mean of the two critics - rather than the minimum - alleviates performance loss here as well. Of note, the alleviation is more pronounced in the same environments where underestimation mitigation helped the most with smaller actors (fig. 4). This suggests that addressing underestimation in SAC can additionally help mitigate the challenges arising from partial observability.
@@
In general, developing a greater understanding between optimization, exploration, expressivity, and estimation accuracy will lead to more robust and reliable reinforcement learning agents. While our work has focused on the case of small actors, the insights provided help strengthen our collective understanding of these learning dynamics. Addressing overestimation in AC methods by taking the minimum of estimated Q values has been a continuing trend - for example, with Deep Deterministic Policy Gradient (DDPG; Lillicrap et al., 2016) being followed by Twin-Delayed DDPG (TD3; Fujimoto et al., 2018). However, our work contributes to findings showing that this approach contributes to underestimation, which warrants further consideration particularly in settings where data collection is more challenging (Ciosek et al., 2019; Li et al., 2023b; He & Hou, 2020). Further, all these considerations are aligned with the continued relevance of the exploration-exploration dilemma (Li et al., 2023b; Sutton & Barto, 2018), which has been explored via Thompson sampling-like techniques (Ishfaq et al., 2025) and through information gain maximization (Sukhija et al., 2025).
@@ 6 Extra results
Figure 8: Decreasing the size of the actor in DrQ decreases validation Q values, as described in fig. 3. We report Q values at the end of training relative to the default baseline. The solid lines represent mean performance, while the shaded area represents the $95\%$ confidence interval, computed across $10$ seeds.
@@
Figure 13: The impact of small actors on a number of metrics related to plasticity as measured on the critics. We also evaluate these metrics when using the mean and max of the two critics, as discussed in section 4.1.
Figure 14: The impact of small actors on a number of metrics related to plasticity as measured on the actor. We also evaluate these metrics when using the mean and max of the two critics, as discussed in section 4.1. We define $o_{\phi}$ on the actor as $o_{\phi}:=\frac{\mathbb{E}_{\mathcal{D}}H}{\mathbb{E}_{\mathcal{D}_{V}}H}$, where $H$ is the entropy of the actor’s action distribution, and $\mathcal{D}_{V}$ is a validation dataset.

### v3

@@ 1 Introduction
The paper is organized as follows: in section˜2, we lay the groundwork and explain our experimental setup. In section˜3, we show the performance effects when naively reducing a smaller actor across a variety of state-based and image-based continuous control tasks, and analyze what could be the cause of performance differences. In section˜4, we focus on interventions that can gain back performance, specifically focusing on bias correction and value function underestimation. Finally, we conclude with discussions and avenues for future work in section˜5.
@@ 2 Preliminaries
In eq.˜2, $V_{\bar{\phi}}$ is the value function computed from $Q_{\phi}$ via $V_{\bar{\phi}}(x)=\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid x)}\left[Q_{\bar{\phi}}(x,a)-\alpha\log\pi_{\theta}(a\mid x)\right]$, where $\bar{\phi}$ are delayed target parameters (Mnih et al., 2015), and $\alpha$ is a learned Lagrange multiplier (we exclude its parameterization for simplicity of exposition). In their practical implementation, Haarnoja et al. (2018b) use two $Q$ value estimates with parameters $\phi_{1}$ and $\phi_{2}$, trained independently, and take their minimum in the update terms in equations 2 and 3, resulting in the following updated losses, with $V_{\bar{\phi}}(x)=\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid x)}\left[\min_{i\in\{1,2\}}Q_{\bar{\phi_{i}}}(x,a)-\alpha\log\pi_{\theta}(a\mid x)\right]$:
@@
It is important to note the interconnectedness of these losses: the actor influences the critic via the (soft) value function $V_{\bar{\phi}}$ used in eqs.˜2 and 4, while the critic influences the actor via $Q_{\phi}$ in eqs.˜3 and 5. Additionally, the actor influences the training dynamics of both given that it is in charge of data collection. Finally, note the use of the TD-error term in eqs.˜2 and 4.
@@ 3 The impact of small actors
In the top row of fig.˜1 we evaluate the impact on performance when reducing the size of these layers and can see a clear degradation in performance across all environments. We additionally measure $o_{\phi}$ on the critics, introduced by Nauman et al. (2024) as a measure of overfitting, defined as $o_{\phi}:=\frac{\mathbb{E}_{\mathcal{D}_{V}}TD_{\phi}}{\mathbb{E}_{\mathcal{D}}TD_{\phi}}$. Here, $\mathcal{D}_{V}$ is a validation dataset of size $11,000$, containing data sampled from a training run with a regular unmodified SAC agent, trained with a different random seed, and $TD_{\phi}$ is the temporal difference error. Higher values of $o_{\phi}$ are indicative of overfitting which, as seen in the middle row of fig.˜1, are inversely correlated with the size of the actor. Finally, we report the fraction of dormant neurons, defined as the proportion of neurons that are 0 for every data point in the validation buffer, where higher levels of dormancy is associated with a loss of plasticity (Sokar et al., 2023; Lyle et al., 2024; Klein et al., 2024). In the bottom row of fig.˜1 we see that the fraction of dormant neurons tends to be inversely correlated with actor size and performance, particularly for the environments where the performance loss is greatest, although to a lesser extent than $o_{\phi}$.
@@ 3.1 Smaller actors collect worse data
Training overparameterized neural networks on small datasets is a common cause for overfitting. Given that the actor is in charge of data collection and reducing its expressivity results in overfit critics, we continue our investigation by evaluating the quality of the data gathered by differently-sized actors. For this, we evaluate training on data collected by separate, and differently-sized, actors. Specifically, we train the regularly-sized actor $\pi_{r}$ with data provided by one of the smaller actors, where the data collection exactly mimics that obtained by the smaller actor during training. This is depicted by the black bars in the top row of fig.˜2, where we can see the performance to be clearly correlated with actor size.
Nikishin et al. (2022) demonstrated the tendency of RL agents to overfit to early experience, affecting their plasticity and downstream performance. It is thus worth considering whether the quality of the training data on an actor is most important in the early stages of training. To evaluate this, in the bottom row of fig.˜2 we explore training the smallest actor ($\pi_{xs}$) on data provided by the $\pi_{r}$ actor, again matching the data collection of the smaller actor. Our analyses here explore using the data from $\pi_{r}$ for only a fraction of training, and then switching to data collected by $\pi_{xs}$ itself. As more data is collected from the bigger actor $\pi_{r}$, the performance of $\pi_{xs}$ generally improves. We note that using all of the data from $\pi_{r}$ (i.e. at 100%) sometimes results in degraded performance; we hypothesize that this may be due to the tandem effect observed by Ostrovski et al. (2021). Overall, we see an improvement in performance in the environments most impacted by reducing the size of the actor for SAC. With DrQ, we do not see a pronounced effect when training the smallest actor on data from a regular-sized actor (see fig.˜9), but similarly, a small trend may be observed for environments with the biggest degradation in performance with reduced actor sizes.
@@ 3.2 Smaller actors result in critic underestimation
The reduction in the quality of data gathered by small actors can possibly be attributed to under-exploration of the state space. This can often be a consequence of an overly-conservative critic which under-estimates values, as well as a low-entropy actor with low diversity in action selection. In fig.˜3 we compare the average critic validation $Q$-values (computed on the same validation dataset) as well as the entropy of the actor’s action distribution $\pi$ of the smaller actors relative to the regularly-sized actor, and confirm that smaller actors result in $Q$-value underestimation, as well as a reduction in entropy during training (see fig.˜11 for the same comparison throughout training). The observed underestimation is interesting, given that it stands in contrast to the over-estimation that’s commonly addressed in many popular algorithms (Hasselt, 2010; Hasselt et al., 2016; Fujimoto et al., 2018).
@@ 4.1 Average and maximal critics
We begin by a simple modification to the original SAC losses to directly address the observed value underestimation. Specifically, we replace the minimization of the two independent $Q$ estimates in equations 4 and 5 with either their mean ($\textrm{avg}(Q_{\phi_{1}},Q_{\phi_{2}})$) or their maximum ($\max(Q_{\phi_{1}},Q_{\phi_{2}}))$. As can be seen in the top and middle rows of fig.˜4 and the top row of fig.˜5, this approach can be quite effective at boosting the performance of small actors in SAC, sometimes even improving over the minimization approach with the regular sized model (e.g. hopper-hop). The bottom row of fig.˜5 confirms that this technique does increase the validation value estimates. As can be seen in fig.˜13, we find that the mean and the max approaches also improve several overfitting and plasticity metrics in the critics, most notably $o_{\phi}$ and the rank of the last hidden layer (Kumar et al., 2021; Nauman et al., 2024). However, they do not appear to have a notable impact on these metrics in the actor (see fig.˜14). The results on the smallest actor on DrQ (bottom row of fig.˜4) display a similar performance trend, although the results are less pronounced. We also observe a corresponding trend with an increase in validation Q values with the mean and max approaches in DrQ in fig.˜10.
@@ 4.2 Critic regularization
Prior work has shown that resetting or regularizing the critic in particular improves plasticity (Ma et al., 2023; Nikishin et al., 2022; Liu et al., 2021) and can help mitigate value-estimation issues, particularly in the case of layer normalization, (Nauman et al., 2024), albeit with overestimation. Given both the increased overfitting observed in the critics, and how much value estimation is affected by smaller actors (see fig.˜3), we investigate whether critic regularization alone can be effective mitigating this impact by applying a number of regularization techniques, focusing on SAC: (a) Layer Normalization (Ba et al., 2016); (b) Spectral Normalization (Miyato et al., 2018); (c) weight decay (van Laarhoven, 2017) with a regularization value of $0.01$ (Li et al., 2023a); (d) L2 distance from initialization (Kumar et al., 2024): with a value of $1\times 10^{-7}$ after tuning on the range $5\times 10^{-8},1\times 10^{-4}]$ in increments of $0.5$ with quadruped-run; and (e) Network resets: resetting neural network layers during training (Nikishin et al., 2022). We apply layer normalization and spectral normalization to the second hidden layer in the critics, and we reset only the output layer of the critics every 5050K steps. Although many of these methods do appear to help with mitigating value under-estimation (bottom row of fig.˜5), they do not appear to help much with performance (top row of fig.˜5 and fig.˜12). For DrQ, we investigate resetting the MLP of the critics (Nikishin et al., 2022) for the smallest actors, and similarly do not see a notable rescue effect (see fig.˜10).
@@ 4.3 Addressing bias in the critic via actor representations
In asymmetric actor-critics methods, imbalances in information received by the critic versus the actor can lead to biased gradients that may negatively impact performance; Baisero & Amato (2022) and Lyu et al. (2022) propose to alleviate this by giving the (limited) information received by the actor as additional input into the critic. In our case, the critics do not receive privileged information over the actor, but we theorize that a similar effect may be occurring within the policy network due to potentially impacted information flow through the smaller actors. We attempt a similar bias correction by concatenating the latent state of the final hidden layer of the actor as input to the final hidden layer of the critics. The latent state is first projected through an untrained neural network layer to a size of $8$ to maintain consistency across actor sizes. As shown in the top row of fig.˜5 and the bottom row of fig.˜12, the bias correction performs similarly to other attempted critic regularization methods in SAC.
@@ 5 Discussion
Real-world problems are often subject to constraints such as latency, model size, and interpretability, which are largely absent in the academic benchmarks where machine learning solutions are developed. As such, it is imperative that we develop the necessary techniques for training reinforcement learning agents under such limitations. The use of small actors can help reduce latency, memory, and inference costs, and can help improve interpretability; these are all practical considerations, as ultimately it is a trained actor which will be deployed for action selection. Our work demonstrates that naïvely shrinking the actor can result in value underestimation, poor data collection, and ultimately degraded performance. We evaluated a number of approaches for mitigating this deterioration and found the most effective to be simply replacing the min operation with a mean or max when combining the values of the two critics (section˜4.1).
It is often necessary to provide the actor with less information than the critic, as was employed by Vasco et al. (2024) to better match the inputs used by humans. In fig.˜6 we explore whether this additional type of limitation on the actor may have a similar effect to what we observe when decreasing the size of the actor. To test this, we zero out two-thirds of the actor inputs in SAC (retaining every third dimension) and find that taking the mean of the two critics - rather than the minimum - alleviates performance loss here as well. Of note, the alleviation is more pronounced in the same environments where underestimation mitigation helped the most with smaller actors (fig.˜4). This suggests that addressing underestimation in SAC can additionally help mitigate the challenges arising from partial observability.
@@
In general, developing a greater understanding between optimization, exploration, expressivity, and estimation accuracy will lead to more robust and reliable reinforcement learning agents. While our work has focused on the case of small actors, the insights provided help strengthen our collective understanding of these learning dynamics. Addressing overestimation in AC methods by taking the minimum of estimated Q values has been a continuing trend - for example, with Deep Deterministic Policy Gradient (DDPG; Lillicrap et al., 2016) being followed by Twin-Delayed DDPG (TD3; Fujimoto et al., 2018). However, our work contributes to findings showing that this approach contributes to underestimation, which warrants further consideration particularly in settings where data collection is more challenging (Ciosek et al., 2019; Li et al., 2023b; He & Hou, 2020). Further, all these considerations are aligned with the continued relevance of the exploration-exploitation dilemma (Li et al., 2023b; Sutton & Barto, 2018), which has been explored via Thompson sampling-like techniques (Ishfaq et al., 2025) and through information gain maximization (Sukhija et al., 2025).
@@ 6 Extra results
Figure 8: Decreasing the size of the actor in DrQ decreases validation Q values, as described in fig.˜3. We report Q values at the end of training relative to the default baseline. The solid lines represent mean performance, while the shaded area represents the $95\%$ confidence interval, computed across $10$ seeds.
@@
Figure 13: The impact of small actors on a number of metrics related to plasticity as measured on the critics. We also evaluate these metrics when using the mean and max of the two critics, as discussed in section˜4.1.
Figure 14: The impact of small actors on a number of metrics related to plasticity as measured on the actor. We also evaluate these metrics when using the mean and max of the two critics, as discussed in section˜4.1. We define $o_{\phi}$ on the actor as $o_{\phi}:=\frac{\mathbb{E}_{\mathcal{D}}H}{\mathbb{E}_{\mathcal{D}_{V}}H}$, where $H$ is the entropy of the actor’s action distribution, and $\mathcal{D}_{V}$ is a validation dataset.